• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jwexler

    Log entry time 08:06:43 on June19,2010

    Entry number 327508

    keyword=Shift summary - Owl 6/19

    0:05 - Interrupting run 4682 to follow Kent's run plan. Starting with large charge asymmetry test. MCC took away beam between stop of production and start of test, waiting for beam.

    0:20 - Entering large charge asymmetry mode. Changed PC setpoint. MCC unable to change RHWP setting, stuck at 2350. Proceed with less-large charge asymmetry.

    1:10 - Finish large charge asymmetry run, run #4685, beginning BPM calibration.

    1:30 - Starting run with BPM cal. run, #4687 with gain search mode set to manual, ioc10 set to 'gains off'.

    2:00 - Resuming production, run plan calling for one run at 50uA, one at 60uA, then leaving the current at 70uA.

    2:10 - Low beam cut is too high. Lowered it, started a new run.

    3:50 - Finished 50uA, 60uA runs.

    4:30 - Compton EPICS died. Phoned an expert. Rebooted and resuming running.

    5:40 - Flipped HWP. Now on slug 38, IHWP OUT.

    7:30 - Tried to do raster scan on target. Spectrometer DAQs stuck. trying to fix at end of shift.

    Runs:
    Junk: 4683, 4684, 4686 4688-90

    Large charge asym: 4685
    BPM Cal.: 4687

    50uA Production: 4691
    60uA Production: 4692
    70uA Production: 4693-6